OpenStudy (anonymous):

Write an equation of the line, in point-slope form, that passes through the two given points. points: (–17, 8), (3, —2)

OpenStudy (anonymous):

to get your slope you have to do y1-y2 /over x1-x2 then point slpe form is y-y2=m(x-x2) (m=slope)